Circular No. 6505

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996bu IN NGC 3631 S. Nakano, Sumoto, Japan, communicates that Reiki Kushida, Yatsug-

atake South Base Observatory, reports her discovery on an unfiltered CCD image taken on Nov. 14.793 UT with a 0.40-m f/5 reflector of an apparent supernova (mag 17.3) located at α = 11^{h}20^{m}59^{s}.18, δ = +53^{o}12′08′′.0 (equinox 2000.0; measurement by Y. Koshida), which is 32′′ west and 117′′ north of the center of NGC 3631. No star at this position is visible on her frames taken on 1996 Feb. 13 and May 9.

P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that C. Hergenrother and W. Brown obtained CCD images of SN 1996bu with the Whipple Observatory’s 1.2-m reflector on Nov. 17.49 UT, yielding V = 17.6 and B = 18.1 and position end figures 59^{s}.30, 08′′.4 for the supernova. Spectra obtained with the Multiple Mirror Telescope by R. Marzke and S. Odewahn on Nov. 18.5 show SN 1996bu to be a peculiar type-II supernova. Narrow Balmer emission lines are seen above a blue continuum; the continuum becomes flat shortward of 500 nm. The widths of the lines are 1400 km/s FWHM after correcting for the spectral resolution; at the base of the Hα emission, the full width increases to 5500 km/s. The Hα emission centroid is redshifted by 1140 km/s, which is close to the velocity of 1156 km/s for NGC 3631. There is no evidence of broad P-Cyg features seen in normal type-II supernovae. Deep absorption features of Na I 590.7-nm (0.4 nm equivalent width) and Ca II 395.7-nm (1.5 nm EW) are seen and may be due to interstellar or circumstellar gas.

LEONID METEORS 1996 Numerous observers have reported observations that suggest a mod-

erate 1996 display of particularly bright Leonid meteors. J. Rendtel and R. Arlt, International Meteor Organization, report no distinct peak but rather a plateau of activity (ZHR > 30; calculated with population index 1.8) between Nov. 17.05 and 17.60, with somewhat higher rates (ZHR ∼ 60) during Nov. 17.2–17.4. J. E. Bortle, Stormville, NY, reports that the display on Nov. 17 UT was similar to the 1961 display of unusually bright meteors. P. Jenniskens (Ames Research Center), P. Brown (University of Western Ontario), and R. Keen (Coal Creek Canyon, CO) also remarked on the bright meteors, Keen noting that all Leonids observed during Nov. 17.32–17.42 left trains that lasted in most cases for about one second (the brightest such meteor having mag −6), and Brown noting seven fireballs in the magnitude range −4 to −8 during Nov. 17.19–17.45.

1996 November 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green